OpenStudy (anonymous):

Solve the problem. Chandra has 5 liters of a 24% solution of sodium hydroxide in a container. What is the amount and concentration of sodium hydroxide solution she must add to this in order to end up with 8 liters of 27% solution? Choose one answer. a. 3 liters of 32% solution b. 3 liters of 30% solution c. 3 liters of 35% solution d. 3 liters of 33% solution

OpenStudy (anonymous):

5*0.24+3*x=8*0.27 x=0.32 so a. is the answer